What is the limiting reactant when 1.50g of magnesium and 1.50g of nitrogen combine according to the equation:
3Mg(s) + N2(g) ⇨ Mg3N2(s)
and how many grams of Mg3N2 are formed?
1
Expert's answer
2020-07-07T14:20:58-0400

3Mg(s) + N2(g) "\implies" Mg3N2(s)

Moles of Mg = 1.5/24

Moles of N2 = 1.5/28

For finding limiting reagent we will divide no. of moles by stoichiometric coefficient in balance reaction and find which value is less that will be the limiting reagent.

So on solving we find that Mg will be the limiting reagent due to less value.


Now after finding reagent apply mole concept.

Mole concept is always based on limiting reagent.

Hence the no. of moles of Mg3N2 will formed is calculated via unitary method.

Unitary method

3 mole Mg will give 1 mole Mg3N2

So 1.5/24 moles will give 1.5/24*3 moles Mg3N2

Hence mass of Mg3N2 = (1.5/72)*100

Since molecular mass of Mg3N22 is 100 gm

Mass of Mg3N2 formed = 2.0833 gm.

This is the answer.

Mass = moles * molecular mass


Stoichiometric coefficient are no. of moles used in a balance reaction.